OpenStudy (anonymous):

Becca's family has 4 gallons of milk. They drink gallon per day.

OpenStudy (leader):

Tell Becca's family that drinking milk is good for them.

OpenStudy (anonymous):

they drunk 5/8 milk evry day*

OpenStudy (anonymous):

drink*

OpenStudy (leader):

every*

OpenStudy (anonymous):

if your not gonna be helpful please go do it somewhere else

OpenStudy (leader):

So if they drink 5/8 gallons everyday and now they have 4 gallons. How many days can they drink milk for?

Mehek (mehek14):

so what is the question

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

at the top

Mehek (mehek14):

ok so 4 * 5/8

OpenStudy (anonymous):

Becca's family has 4 gallons of milk. They drink gallon per day. How long will their milk last?

OpenStudy (anonymous):

5/8*

OpenStudy (anonymous):

how many days would it take to finish the milk?

OpenStudy (anonymous):

yea

OpenStudy (leader):

\(\color{blue}{\text{Originally Posted by}}\) @Leader So if they drink 5/8 gallons everyday and now they have 4 gallons. How many days can they drink milk for? \(\color{blue}{\text{End of Quote}}\)

Mehek (mehek14):



OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

8gal * 1day/(5/8) gal = ?

OpenStudy (anonymous):

4gal i mean

Mehek (mehek14):

4 / (5/8) = 4 * 8/5 = 32/5 = 6 2/5

Mehek (mehek14):